# Fixed Points of the Josephus Function via Fractional Base Expansions

### v2

@@ 3 Concluding Remarks
In this paper, we studied the fixed point sequence of the Josephus function $J_{3}$ and showed how its structure can be better understood through number theory and a non-integer base representation. First, we used the Chinese Remainder Theorem to describe each fixed point as the solution of a system of two linear congruences, exploiting the fact that the moduli involved are coprime. Then, we explored a numerical pattern that appears when the fixed points are expressed in base $3/2$ and provided a recursive formula for computing the digits of their expansions.
This connection offers a new and simple way to describe the fixed points of $J_{3}$ and opens the door to further analysis using tools from fractional number systems in this setting. Overall, the use of base $3/2$ provides a helpful perspective for understanding the behavior of the fixed points and suggests interesting directions for future research. One such direction is to investigate whether the fixed points of $J_{4}$ admit a recursive formula in base $4/3$, potentially revealing patterns similar to those found for $J_{3}$ and $J_{2}$. However, $J_{4}$ has two distinct pure high extremal points, which may require at least two indices to describe the number of pure high extremal points between consecutive fixed points.
Another open and promising direction is to determine whether the Josephus function itself, namely $J_{3}(n)$, can be computed directly using base $3/2$ representations. In this case, the expression for $J_{3}(n)$ given in (9) and (10) of [2] involves two indices, which complicates the derivation of a recursion for the base $3/2$ expansion of $J_{3}(n)$.
Acknowledgement. The first author would like to acknowledge the support provided by the National Science Foundation (NSF) grant # DMS - 2307328. The authors are grateful to the anonymous referees for their valuable comments and suggestions, which helped improve the quality of the paper.

### v3

@@ 3 Concluding Remarks
In this paper, we studied the fixed point sequence of the Josephus function $J_{3}$ and showed how its structure can be better understood through number theory and a fractional base representation. First, we used the Chinese Remainder Theorem to describe each fixed point as the solution of a system of two linear congruences, exploiting the fact that the moduli involved are coprime. Then, we identified a numerical pattern in the base $3/2$ expansions of the fixed points and provided a recursive formula for computing the digits of these expansions.
This connection offers a new and transparent way to describe the fixed points of $J_{3}$ and opens the door to further analysis using tools from the theory of fractional base representations. The use of base $3/2$ provides a natural perspective for understanding the structure of the fixed points and suggests several directions for future research.
One such direction is to investigate whether the fixed points of $J_{4}$ admit a recursive digit formula in base $4/3$, potentially revealing patterns analogous to those found for $J_{3}$ and $J_{2}$. However, $J_{4}$ has two distinct types of pure high extremal points, which may require at least two indices to track the number of pure high extremal points between consecutive fixed points.
Another open direction is to determine whether $J_{3}(n)$ itself can be computed directly from the base $3/2$ expansion of $n$. The expression for $J_{3}(n)$ given in equations (9) and (10) of [2] involves two indices, which complicates the derivation of a recursion for the base $3/2$ expansion of $J_{3}(n)$.
Acknowledgements. The first author acknowledges the support of the National Science Foundation (NSF) grant #DMS-2307328. The authors are grateful to the anonymous referees for their valuable comments and suggestions, which improved the quality of the paper.